Let A consist of all 52 cards in an ordinary deck of playing cards. Suppose that


this deck is shuffled and a hand of five cards is dealt. A list of cards in this hand,


in the order in which they were dealt, is a permutation of A taken five at a tim

1
Expert's answer
2022-01-24T15:37:03-0500

Solution;

The number of ways of 5 cards drawn at a time is a permutation of;

"(\begin{array}{cc}
 52 \\
 5
\end{array})" ="\frac{52!}{5!(52-5)!}" "=\frac{52×51×50×49×48×46!}{5!×47!}"

"=\frac{311875200}{120}=2598960"